Cash Received from Customers—Direct Method Sales reported on the income statement were $541,000. The accounts receivable balance decreased $45,000 over the year. Determine the amount of cash received from customers.
Answer #1

Let beginning AR=$100,000

Hence ending AR=(100,000-45000)=$55000

Ending AR=Beginnng AR+Sales-Cash collections

Cash collections=(100,000+541000-55000)

=$586,000.